# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v3

@@ Abstract
Figure 1: We provide for the first time 6D pose annotations in the form of a per-image object pose distribution. Current annotations in BOP [17] datasets are given as a single pose, shown here as a circle in the SO(3) representations. T-LESS [13] also provides a symmetry pattern leaving the 3D object model invariant, from which a distribution can be computed (the colored points in the SO(3) representations). Such distribution however does not cover many cases [30]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3). When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 modes. When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose. Our method annotates the T-LESS dataset with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly. We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge. Such ground truth distributions also become a key asset when it comes to evaluating pose distribution estimation methods [9, 19]. With appropriate metrics, we demonstrate the first quantitative evaluation of pose distribution methods on real images, as an extension to single pose methods.
@@ 1 Introduction
Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explains the observed image. It is a key task in many application domains, such as robotics (e.g. grasping or manipulation), augmented reality, industrial quality control, etc. In such contexts, datasets of images with ground-truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches. The annotation accuracy of reference benchmarks, such as BOP [43] for 6D pose estimation, is thus crucial since it influences the research directions.
However, annotating images of objects with their 6D poses is a complex task. This is especially true when considered objects include symmetrical parts. Indeed, as illustrated in Figure 1, an image does not necessarily correspond to a single pose solution depending on the viewpoint and occlusions. While symmetrical objects naturally imply multiple solutions, non-symmetrical objects can also yield multiple solutions in case of partial occlusions. Current reference benchmarks [43] ignore these occlusion-induced symmetries, resulting in an imperfect evaluation of the methods. As we will show, considering better ground truth distributions results in a significant change of ranking of current pose estimation methods.
In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. Its genericity is demonstrated on T-LESS [13] and YCB-V [48].
@@
Moreover, we also consider the few methods that already predict multi-modal pose distributions [9, 19]. Because of the absence of proper annotations, these methods have been evaluated quantitatively only on synthetic images and only qualitatively on real images. Thanks to our annotations, we are able to provide their first evaluation on real data.
@@ 2 Related Work
While the state of the art mostly focused on estimating a single pose [6, 39, 15, 24, 42, 45, 8, 41], some recent works focus on estimating the complete set of solutions [30, 37, 34, 23, 31, 18, 9, 19]. As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are then required, even for methods that return a single pose per image.
@@
To our knowledge, no method has been proposed to determine this ground-truth transformation set per image. Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set, usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g. surface deviation tolerance of 1.5cm or 10% of the object diameter [16]), and cannot be used to evaluate properly multi-modal pose distribution estimates like [9, 19].
@@
Performance evaluation usually differs depending on whether the method is used to output a single pose or a distribution.
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [12]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 48, 44]. In practice, the most commonly used accuracy measures are :
@@
    Maximum Symmetry-Aware Projection Distance (MSPD) [16] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management, and provides an error useful for augmented reality.
@@
[…]
@@ 3 Method
Consider a 3D object model $M$ observed from a pose $P_{\text{GT}}$. We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are indistinguishable from the one generated by $P_{\text{GT}}$. Depending on the object geometry $M$ and the representation type, the set ${\cal T}$ can be limited to the identity transform or can include multiple transformations in addition to the identity.
@@
Consequently, to estimate ${\cal T}(I)$, our solution compute first ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (section 3.3). As illustrated in Figure 2, to achieve this estimation efficiently, our solution decompose the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first step of pre-computation (section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that compute ${\cal T}({\cal V}(M))$ (section 3.2).
@@ 3.3 Refining $\epsilon$-sym: visual occlusions consideration
[…]
@@ 5 Experiments
In this section, after evaluating the performances of our annotation method (Section 5.2), we present and discuss the impact of our ground-truth annotations and evaluation metrics onto the evaluation of state-of-the-art solutions for both single pose estimation (Section 5.3) and multi-modal pose distribution estimation (Section 5.4).
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and HB [22] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths. HOPE [44] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
@@ 5.1 Implementation details
[…]
Surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
5.2 Qualitative validation of the method
We illustrate qualitatively the results obtained on some images of the T-LESS dataset in Figure 3. We can observe the coherence between the distribution and the visibility of the disambiguating elements. Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
Figure 3: Visualizations of our ground truth. We display SE(3) ground truth distributions for scene 16 of T-LESS [13]. Circle on orientation diagram represents the unique ground truth pose provided as input to our method. Colors are used to link object instances to their distributions. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [17] submissions have become publicly available to allow in-depth analysis. This experiment reprocess these results to compare the performance evaluations of state-of-the-art methods for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [9, 19] for which we kept only the highest confidence mode. Since these methods are not part of the BOP challenge, we retrained them.
@@ Results.
It is also interesting to note that the ranking order of variants v1.2, v1.3 and v1.4 of GDRNPP RGB-D (based on different object detectors, training augmentations, and refinement methods) is changed when evaluated with our ground truth. It underlines that inaccurate ground truth annotation can induce erroneous conclusions in ablation studies.
@@
We evaluated SpyroPose [9] and LiePose-Diffusion [19] (either PBR + real images and PBR only to compare with SpyroPose) since they provide 6D pose distributions, unlike [34, 18] that provide distributions only over the rotation. To show genericity, we also evaluated the top 3 RGB methods of single pose estimation task (with image-wise annotations). Their distribution is considered as a Dirac distribution on the single estimate. Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30, as illustrated in supplementary material).
@@ 6 Limitations
Our annotation method relies on a geometric analysis of the scene. It is possible that the sensor resolution, sensor noise, field of view, or motion blur may also affect the visibility of the disambiguating parts. Considering these effects would improve the ground truth even further.
@@ 7 Conclusion
For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, completed by a per-object symmetry pattern to transform this unique pose into a distribution. We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting distribution ground truth and in the resulting performances evaluations.
We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries. We illustrated that the resulting ground truth is more accurate. Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically. We also introduced metrics to evaluate methods that estimate a pose distribution and provided their first evaluation on real data.
We believe that these results should encourage the 3D vision community to adopt a new outlook on these methods and how well they work in practice, and should guide future research. To facilitate development, we will make our annotations and code publicly available.
@@ 9 Sort(H)
Listing 2: Image annotation, pseudo-code of equation (4) from section 3.2.
@@
Listing 3: Image annotation depth post-processing, pseudo-code of equation (4) from section 3.3, based on VSD implementation from BOP toolkit.
@@ 2 Analysis of our BOP-Distrib Annotations
2.1 Accuracy of the annotations
Figure 4 illustrate renderings of all poses of a symmetries pattern.
Figure 4: Variation between poses in a single mode symmetries pattern. We display the renderings of the symmetries pattern of object 2 in image 7, scene 1. Although the SO(3) visualization may show spread points, these rotations result in similar poses.
@@ 2.2 False visible points and false occluded points detected by the pruning stage
[…]
@@ 2.4 T-LESS [13] Annotation Details
[…]
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.

### v4

@@ Abstract
Figure 1: We provide for the first time 6D pose annotations in the form of a per-image object pose distribution. Current annotations in BOP [21] datasets are given as a single pose, shown here as a circle in the SO(3) representations. BOP also provides a symmetry pattern per object, from which a distribution can be computed (the colored points in SO(3)). Such distribution however does not cover many cases [35]: In this example, when only the core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3). When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 modes. When the hole is visible (Case 3), the pose distribution should be concentrated around one non-ambiguous pose. Our method annotates scenes with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly. We show that considering these distributions for evaluation results in a significant change of ranking for the BOP challenge. Such ground truth distributions also become a key asset when it comes to evaluating pose distribution estimation methods [13, 23]. With appropriate metrics, we demonstrate the first quantitative evaluation of pose distribution methods on real images, as an extension to single pose methods.
@@ 1 Introduction
Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explains the observed image. It is a key task in many application domains, such as robotics (e.g. grasping or manipulation), augmented reality, industrial quality control, etc. In such contexts, datasets of images with ground truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches. The annotation accuracy of reference benchmarks, such as BOP [47] for 6D pose estimation, is thus crucial since it influences the research directions.
However, annotating images of objects with their 6D poses is a complex task. This is especially true when considered objects include symmetrical parts. Indeed, as illustrated in Figure 1, an image does not necessarily correspond to a single pose solution depending on the viewpoint and occlusions. While symmetrical objects naturally imply multiple solutions, non-symmetrical objects can also yield multiple solutions in case of partial occlusions. Current reference benchmarks [47] ignore these occlusion-induced symmetries, resulting in an imperfect evaluation of the methods. Considering better ground truth distributions results in a significant change of ranking of pose estimation methods.
In this paper, we propose a method that automatically annotates pose distributions for images that take into account ambiguities due to object symmetries but also to partial occlusions. Our method exploits the current annotations, the ground truth visibility masks of the objects, and their 3D models, to produce a non-parametric distribution representation (see Supp. Mat. Section 2.2). Its genericity is demonstrated on T-LESS [17] and YCB-V [54].
@@
Moreover, we also consider the few methods that already predict multi-modal pose distributions [13, 23].
Because of the absence of proper annotations, these methods have been evaluated quantitatively only on synthetic images and only qualitatively on real images. Thanks to our annotations, we are able to provide their first evaluation on real data.
@@ 2 Related Work
While the state of the art mostly focused on estimating a single pose [10, 44, 16, 28, 48, 51, 12, 46], some recent works focus on estimating the complete set of solutions [35, 42, 39, 27, 36, 22, 13, 23, 11, 7, 4, 55, 30, 50]. As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are then required, even for methods that return a single pose per image.
@@
To our knowledge, no method has been proposed to determine this ground truth transformation set per image. Instead, the current gold standard in 6D pose estimation benchmarks still consists in approximating this per-image transformation set with a unique global transformation set, usually computed with the same method than fully symmetric object with large enough tolerance to ignore small disambiguating elements (e.g. surface deviation tolerance of 1.5cm or 10% of the object diameter [20]), and cannot be used to evaluate properly multi-modal pose distribution estimates like [13, 23].
@@
Performance evaluation usually differs if the method outputs a single pose or a distribution.
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [18]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground truths due to a symmetrically tagged object, the accuracy corresponds to the minimal error with respect to the set of solutions [18, 54, 49]. In practice, the most commonly used accuracy measures are :
@@
    Maximum Symmetry-Aware Projection Distance (MSPD) [20] measures the maximum reprojection error between projections of the ground truth and estimated surface points, with symmetry management.
@@
Figure 2: Method overview. From a symmetry candidate set, we pre-compute the object per-vertex $\epsilon\text{-sym}$. Then for a given scene, we compute the vertices visibility ($\checkmark$ and illustrate respectively if the visibility test passed or not for the vertex) and perform a robust intersection between their $\epsilon\text{-sym}$. This intersection is then pruned with a depth comparison and the result constitutes the symmetries pattern of this object instance for this image. When multiplied by the ground truth, we obtain the SE(3) distribution of the object instance.
@@ 3 Method
Consider a 3D object model $M$ observed from a pose $P_{\text{GT}}$. We call symmetry-pattern the set ${\cal T}=\{T_{i}\in{\text{SE}}(3)\}$ of rigid transformations that, once combined with $P_{\text{GT}}$, generate representations that are geometrically indistinguishable from the one generated by $P_{\text{GT}}$. Depending on the object geometry $M$ and the representation type, the set $\cal T$} can be limited to the identity transform or can include multiple transformations in addition to the identity.
@@
Consequently, to estimate ${\cal T}(I)$, our solution computes ${\cal T}({\cal V}(M))$ then prune the pose visually-inconsistent (Section 3.3). As illustrated in Figure 2, to achieve this estimation efficiently, our solution decomposes the estimation of ${\cal T}({\cal V}(M))$ in a two step process: a first pre-computation (Section 3.1) independent of ${\cal V}(M)$ and executed only once, and a second step that computes ${\cal T}({\cal V}(M))$ (Section 3.2).
@@ 3.3 Refining $\epsilon$-sym: visual occlusions consideration
Since $\epsilon\text{-sym}$ does not consider visual occlusions, ${\cal T}({\cal V}(M))$ may contain a small set of pose that could induce more visible part of $M$ than ${\cal V}(M)$ (see Supp. Mat. Section 2.1).
We follow the VSD metric [18] and determine if a pose $T_{i}$ should be pruned by rendering the depth image $D_{T_{i}}$. For annotation purposes, the minimal disambiguation element size to reject a candidate pose should be the same for all objects. So we remove VSD dependence to object size and count pixels with a depth deviation to ground truth greater than $\delta$ (mm). If the number of pixels exceeds $\tau_{pix}$, the pose is pruned.
@@ 5 Experiments
In this section, after evaluating the performances of our annotation method (Section 5.2), we present and discuss the impact of our ground truth annotations and evaluation metrics onto the evaluation of state-of-the-art solutions for both single pose estimation (Section 5.3) and multi-modal pose distribution estimation (Section 5.4).
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [21], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground truth annotations: ITODD [9] and HB [26] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground truths. HOPE [49], Omni6D [56] and IC-BIN [8] have symmetrical shapes, but texture information disambiguates them.
@@ 5.1 Implementation details
In our experiments, we sample uniformly surface points from the CAD models with a resolution of 0.5mm. For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 1mm. Following BOP, continuous symmetries are discretized such as the furthest vertex from the axis of symmetry travels not more than 1% of the object diameter.
Surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating size of roughly $\times2.5$mm2, $\delta=5$mm and $\tau_{pix}=30$pix.
5.2 Validation of the method
Figure 3 illustrates the coherence between the distribution and the visibility of the disambiguating elements. The ground truth translation is always reduced to a single 3D point as T-LESS exhibit no translational ambiguity.
Regarding quantitative evaluation of the symmetry pattern accuracy, Table 1 reports the max deviation error (in mm) between the depth image rendered from the GT pose $P_{GT}$ and the ones rendered from the combination of $P_{GT}$ with any pose of the symmetry pattern. We observe that the rendered depths are really close.
$\mu_{\lvert\delta\rvert}$ ($m$) $\text{med}_{\lvert\delta\rvert}$ ($m$) max|δ| (mm) maxVSD maxPrune (pix)
0.218 5.85e-3 0.497 0.061 30
Table 1: Quantitative accuracy analysis of our new pose distribution annotations on T-LESS. We report the differences between rendered depths at the ground truth pose and all the points of the corresponding distribution. $\mu_{\lvert\delta\rvert}$, $\text{med}_{\lvert\delta\rvert}$ and max|δ| are the mean, median and max of the absolute differences, while maxVSD and maxPrune are max errors of VSD@0.05 and of our pruning.
Figure 3: Visualizations of our ground truth. We display SE(3) ground truth distributions for scene 1 of T-LESS [17]. Circle on orientation diagram represents the unique ground truth pose provided as input to our method. Colors link objects to their distributions.
@@ 5.3 Single Pose Estimation Evaluation
Since May 2023, raw results of BOP Challenge [21] submissions have become publicly available to allow in-depth analysis. This experiment reprocess these results to compare the performance evaluations of state-of-the-art methods for both the original 6D annotations and our annotations. We considered top contenders of the BOP challenge for the pose estimation and pose estimation for unseen objects tasks. We also included two methods for the multi-modal pose distribution estimation task [13, 23] for which we kept only the highest confidence mode.
@@ Results.
We evaluated SpyroPose [13] and LiePose-Diffusion [23] (either PBR + real images and PBR only to compare with SpyroPose) since they provide 6D pose distributions, unlike [39, 22] that provide distributions only over the rotation. To show genericity, we also evaluated the top 3 RGB methods of single pose estimation task (with image-wise annotations). Their distribution is considered as a Dirac distribution on the single estimate. Results are given on the complete dataset, and the subset of objects with cylindrical symmetry (1-4, 13-18, 24 and 30).
@@ 6 Limitations
Our annotation method relies on geometric analysis. Sensor resolution, sensor noise, field of view or motion blur may also affect disambiguating parts visibility. Considering these effects would improve the ground truth even further.
@@ 7 Conclusion
For simplicity, most of the 6D pose estimation benchmarks rely on a single 6D pose annotation per image, completed by a per-object symmetry pattern to transform this unique pose into a distribution. We argued that ignoring the per-image nature of the symmetry pattern is prone to bias in the resulting ground truth and performances evaluations.
We then proposed a method to annotate 6D pose distribution with a per-image analysis of the object symmetries. We illustrated that the resulting ground truth is more accurate. Moreover, when using this ground truth to re-evaluate current state-of-the-art methods, we showed that the ranking of these methods changes drastically.
We also introduced metrics to evaluate methods that estimate a pose distribution and provided their first evaluation on real data. Such pose distribution qualification is crucial for downstream tasks. Indeed, access to multiple accurate pose solutions allows the selection of the best one for the task (e.g., considering obstacles) or, when the object is known to be asymmetrical, helps determine if the task can’t be completed from the current viewpoint due to occlusion of the elements that break the symmetry and select the next best view.
Acknowledgment
This work was partly funded by the European Union’s Horizon Europe Research and Innovation Program under Grant 101070227 (CONVINCE) and partly funded by the European Union’s Horizon Europe Research and Innovation program under grant agreement nº 101135708 (JARVIS Project). The authors would like to thank Rasmus Laurvig Haugaard for providing SpyroPose [13] implementation for baseline evaluation.
@@ 4 SampledModel = uniform_resampling(ModelCAD, resolution)
Listing 1: Offline $\epsilon\text{-sym}$ pre-computation with all vertices visible, pseudo-code of Equation 3 from Section 3.1.
@@ 9 Sort(H)
Listing 2: Image annotation, pseudo-code of Equation 4 from Section 3.2.
@@
Listing 3: Image annotation depth post-processing, pseudo-code of Equation 4 from Section 3.3, based on VSD implementation from BOP toolkit.
@@ 2.1 False visible points and false occluded points detected by the pruning stage
Figures 4 and 5 present the results of our pruning procedure. Each $\epsilon\text{-sym}$ mode is rendered to be compared to the ground truth pose depth rendering. The pixel with deviation greater than $\delta$ are counted, and if they too numerous (more than $\tau_{pix}$), the mode is pruned, as in figure 5.
Figure 4: Depth deviation post-processing analysis. For a given image, we display the depth renderings of the ground truth pose and of one $\epsilon\text{-sym}$ mode (1 here). They align well.
Figure 5: Depth deviation post-processing analysis. For a given image, we display the depth renderings of the ground truth pose and of one $\epsilon\text{-sym}$ mode (12). Mode 12 generates several falsely occluded pixels (where the hole should be) and falsely visible pixels (where the hole is but shouldn’t be). Mode 12 is rejected by or pruning stage.
2.2 Choice of Distribution Representation
Unlike approaches based on Bingham distributions [7, 4, 11], implicit representations [39, 22], Wigner harmonics [30], matrix Fisher distributions [55] or continuous symmetry groups [2], we do not have a continuous distribution representation. We represent the ground truth pose distributions in a non-parametric way with a set samples of it. Representing a distribution with a large set of samples is common practice and has the advantage of being general: it can represent distributions with no clear analytical form which happen in case of visual ambiguities beyond symmetries, i.e., it can represent distributions with no clear analytical form which happen in case of visual ambiguities beyond symmetries. Moreover, a set of samples permits an efficient performance evaluation.
@@ 2.4 T-LESS [17] Annotation Details
In our experiments, we sample surface points from the CAD models with a resolution of 0.5mm. For the pre-computation of elementary symmetries patterns for a given object, we use the per-object symmetries pattern given by the BOP challenge [21] as the initial symmetry candidates, with a tolerance factor $\epsilon\text{-sym}$ set to 1mm.
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly $\times2.5$mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.